[figure: 9.7 cm]

INELASTIC NEUTRON SCATTERING AROUND THE MAGNETIC BRILLOUIN ZONE CENTRE AND AT THE ZONE

BOUNDARY IN Mn(18% Cu) ALLOY

J. Jankowska-Kisielińska^{a} K. Mikke^{a} and B. Hennion^{b} aInstitute of Atomic Energy , Świerk, Poland bLaboratoire Léon Brillouin, CEN Saclay, France

The temperature dependence of magnetic excitations was studied in the Mn(18% Cu) 70% - single domain sample by inelastic neutron scattering. We observed the magnetic excitations below T_{N} around the point on the MBZ (magnetic Brillouin zone) boundary, equivalent crystallographically to the MBZ centre (”unfulfilled” MBZ centre) and we were able to separate these excitations from the spin waves observed around the MBZ centre. The intensity of neutrons scattered on these excitations decreases with decreasing temperature below T_{N}, in a different way than that predicted by the population factor for spin waves. They persist even at 300K, what is far below T_{N}. We interpret them as dispersionless excitations similar to those observed above the critical temperature.

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : K. Mikke

Address for correspondence : Institute of Atomic Energy Świerk, 05 400 Otwock Poland

Email address : e06kaz@cyf.gov.pl